# Tracing Multilingual Representations in LLMs with Cross-Layer Transcoders

### v1

@@ Abstract
Multilingual Large Language Models (LLMs) can process many languages, yet how they internally represent this diversity remains unclear. Do they form shared multilingual representations with language-specific decoding, and if so, why does performance still favor the dominant training language? To address this, we train a series of LLMs on different mixtures of multilingual data and analyze their internal mechanisms using cross-layer transcoders (CLT) and attribution graphs. Our results provide strong evidence for pivot language representations: the model employs nearly identical representations across languages, while language-specific decoding emerges in later layers. Attribution analyses reveal that decoding relies in part on a small set of high-frequency language features in the final layers, which linearly read out language identity from the first layers in the model. By intervening on these features, we can suppress one language and substitute another in the model’s outputs. Finally, we study how the dominant training language influences these mechanisms across attribution graphs and decoding pathways. We argue that understanding this pivot-language mechanism is crucial for improving multilingual alignment in LLMs.
Tracing Multilingual Representations in LLMs with
Cross-Layer Transcoders
Abir Harrasse1,2thanks: Equal contribution. First author order determined by coin flip. The first author’s work was done during her internship at the Max Planck Institute for Intelligent Systems., Florent Draye1111Models: https://huggingface.co/collections/CausalNLP/multilingual-tinystories-6862b6562414eb84d183f82a and https://huggingface.co/collections/CausalNLP/multilingual-gpt2-models-684ad70e5fb3c84962306af3, Zhijing Jin1,3,4, Bernhard Schölkopf1,5 1Max Planck Institute for Intelligent Systems, Tübingen, Germany 2Mohammed VI Polytechnic University, Morocco 3University of Toronto, Canada  4Vector Institute, Canada 5ELLIS Institute, Tübingen, Germany abir.harrasse@emines.um6p.ma fdraye@tue.mpg.de zjin@cs.toronto.edu
Figure 1: Overview of our study. We analyze the multilingual behavior of LLMs using attribution graphs derived from cross-layer transcoders (CLTs)Ameisen et al. (2025). These graphs provide fine-grained feature representations and capture their interactions more effectively than methods used in prior studies, such as Logit Lens.
@@ 1 Introduction
The inclusion of diverse languages in the large-scale training data of large language models (LLMs) Grattafiori et al. (2024); Team et al. (2024); Brown et al. (2020); Chowdhery et al. (2022) has led to remarkable multilingual capabilities Shi et al. (2022); Conneau et al. (2020a); Workshop et al. (2023) However, the underlying mechanisms driving their multilingual behavior remain poorly understood. For example, do LLMs form shared multilingual representations in their internal layers, so, why does performance still favor the dominant training language, such as English? Prior mechanistic investigations have yielded conflicting findings about multilingual processing. While some work maps inputs into English-aligned latent spaces and suggests that LLMs process non-English inputs through English-biased representations Wendler et al. (2024); Schut et al. (2025), other studies find evidence of shared cross-lingual grammatical structures encoded in common feature directions Brinkmann et al. (2025). Moreover, all existing approaches rely primarily on static probing methods that cannot capture the dynamic transformation of linguistic information across model layers.
To address the aforementioned limitations, we derive attribution graphs from our self-trained cross-layer transcoders Ameisen et al. (2025), a recent advance in mechanistic interpretability, which can track feature interactions across layers and provide the first mechanistic account of how LLMs transition between shared and language-specific processing modes. Our results suggest that all languages have a shared representation in LLMs’ internal layers, which we call as a “pivot language.” We also discover language-specific decoding in layer layers to handle their multilingual output generation.
Depending on our CLT-driven insights, we conduct two causal studies. First, we are able to change the output language by intervening on a small set of high-frequency language features in the final layers decoding from the pivot language to the target language. Further, we investigate the causal effect of training data composition by varying English-to-multilingual ratios. We analyze our re-trained models on different language ratios, and reveal how language distribution during pretraining shapes internal multilingual representations.
Our contributions are as follows:
@@
    We present comprehensive evidence for a pivot language in multilingual LLMs: models form shared multilingual spaces and use consistent mechanisms to process equivalent queries across languages. This holds across training mixtures, suggesting that pivot languages emerge independently of English dominance, which challenges the view that “models think in English”.
@@
    We release eight multilingual models [^1] Models: https://huggingface.co/collections/CausalNLP/multilingual-tinystories-6862b6562414eb84d183f82a and https://huggingface.co/collections/CausalNLP/multilingual-gpt2-models-684ad70e5fb3c84962306af3 (two sizes × four data mixtures) and their corresponding cross-layer transcoders (CLTs) [^2] CLTs: https://huggingface.co/flodraye, trained on balanced language distributions to support cross-lingual analysis and transfer.
@@ 2.1 Multilingual Language Models
Multilingual language models have evolved from early architectures like mBERT Devlin et al. (2019) to sophisticated models such as XLM-R Conneau et al. (2020b), LLaMA Touvron et al. (2023), and Gemma Team et al. (2024). Despite their impressive capabilities, systematic performance disparities persist across languages Pires et al. (2019); Hu et al. (2020). English consistently outperforms other languages, particularly for low-resource languages and complex reasoning tasks Lauscher et al. (2020); Shi et al. (2022). Benchmarks like XTREME Hu et al. (2020) and XNLI Conneau et al. (2020b) have documented these gaps across language families. The English-dominant training data composition has been hypothesized as a primary factor Bender et al. (2021), though the underlying mechanisms driving these disparities remain poorly understood.
@@ 2.2 Internal Representations and the Pivot Hypothesis
Recent investigations have revealed that multilingual LLMs process non-English inputs through English-aligned latent representations. Wendler et al. Wendler et al. (2024) used logit lens analysis to show that intermediate layers predict English tokens before target-language tokens. Similarly, Schut et al. Schut et al. (2025) demonstrated systematic bias toward English-like representations regardless of input language. These findings suggest models internally translate to English, perform reasoning, then translate back, a mechanism that has been leveraged through embedding alignment Lample and Conneau (2019); Artetxe et al. (2018), translation-based fine-tuning Zhu et al. (2024), and romanization strategies for non-Latin scripts Saji et al. (2025).
Prompting strategies also exploit this English-pivot behavior, including direct prompt translation Shi et al. (2022); Ahuja et al. (2023); Etxaniz et al. (2023) and English chain-of-thought reasoning Wei et al. (2022); Vatsal et al. (2025). While some grammatical features exhibit universal representations across languages Brinkmann et al. (2025), lexical and semantic processing appears more language-specific Chi et al. (2020).
These observations align with the recently proposed “Platonic hypothesis” Huh et al. (2024), which suggests that models converge toward abstract, unified representations capturing conceptual essences beyond surface forms. The pivot-language behavior, where models develop language-agnostic representations before language-specific decoding, provides empirical support for this hypothesis in multilingual settings.
@@ 2.3 Mechanistic Interpretability in LLMs
Mechanistic interpretability seeks to understand transformer internals through circuit analysis Olah et al. (2020); Elhage et al. (2021). Early work identified specific circuits for behaviors like indirect object identification Wang et al. (2022) and induction heads Olsson et al. (2022). Sparse Autoencoders (SAEs) decompose representations into interpretable features Cunningham et al. (2023); Templeton et al. (2024), while transcoders directly model MLP outputs rather than autoencoding activations Dunefsky et al. (2024).
Cross-Layer Transcoders (CLTs) advance this approach by assigning distinct decoder matrices per downstream layer, significantly simplifying feature graphs and enabling tractable circuit analysis Ameisen et al. (2025); Hanna et al. (2025). However, CLT applications remain limited—primarily Anthropic’s foundational work Lindsey et al. (2025b), a “greater-than” mechanism study Merullo et al. (2025), and recent open-source extensions Lindsey et al. (2025a). This scarcity reflects the considerable challenges in training CLTs at scale.
For multilingual models, mechanistic analyses are particularly sparse. Beyond Wendler et al.’s logit lens study Wendler et al. (2024) and Anthropic’s case study Lindsey et al. (2025b), our work represents the first comprehensive CLT-based investigation of multilingual processing. Crucially, ensuring fair cross-language comparisons requires training CLTs on balanced multilingual distributions to avoid feature bias toward dominant languages.
@@ 3.1 Mixture of multilingual LLMs
We study the multilingual behavior of LLMs across five languages: English, German, French, Arabic, and Chinese. To investigate how training data composition affects internal representations, we train four GPT-2 style models (12 layers, $\sim$177.6M parameters) on a realistic web mixture (OpenWebText + FineWeb2), and four tiny-stories style models (4 layers and $\sim$68.5M parameters) on a controlled narrative dataset (TinyStories translated into all five languages). Finally, we include an off-the-shelf LLaMA 1B model for comparison.
@@ 3.2 Cross-Layer Transcoders (CLTs)
To probe internal multilingual mechanisms, we employ Cross-Layer Transcoders (CLTs), which map activations between layers to reveal how semantic representations evolve during processing. A cross-layer transcoder consists of neurons (features) divided into encoder and decoder components.
@@
Following Anthropic guidelines Ameisen et al. (2025), the final training objective is
@@ 3.3 Graph attribution
$a_{\ell,k,n}^{\ell^{\prime},k^{\prime},n^{\prime}}=\sum_{\ell\leq s\leq\ell^{\prime}}f_{k,n}^{\ell\to s}\;J_{\ell,k}^{\ell^{\prime},k^{\prime}}\;g_{k^{\prime},n^{\prime}}^{\ell^{\prime}},$ (4)
where $f_{k,n}^{\ell\to s}$ denotes the vector for feature $n$ in the decoder matrix projecting from layer $\ell$ to $s$, $J_{\ell,k}^{\ell^{\prime},k^{\prime}}$ is the Jacobian between the MLP output at $(\ell,k)$ and the MLP input at $(\ell^{\prime},k^{\prime})$, computed during a forward pass where nonlinearities (normalization layers, attention computations, and MLPs) are frozen using stop-gradient operations, and $g_{k^{\prime},n^{\prime}}^{\ell^{\prime}}$ is the corresponding encoder feature at layer $\ell^{\prime}$ and position $k^{\prime}$. The graph is then pruned to retain only features that cumulatively account for $80\%$ of the effect on the final logit, and edges that cumulatively account for $95\%$ of the edge effect on the final logit. We use the circuit-tracer library Hanna et al. (2025).
@@ 3.4 Multilingual Score
For each CLT feature $f$, we compute the multilingual scores on both the top 100 most activated sequences for that feature and on a general sample of 600k sequences (each of length 16). If a feature is active for at least one token in a sequence, it is considered active for that sequence. Since each sequence corresponds to a specific language, this allows us to compute a normalized activation distribution:
@@
which measures multilinguality: low $H(f)$ indicates language-specific features, while high $H(f)$ indicates multilingual features.
@@ 4 Results
    Do models form shared multilingual representations or a pivot language—across data mixtures?
@@
    How do models decode language identity and route information to the correct output language?
@@
    Why do models still fail in non-English settings despite shared representations?
@@ 4.1.1 Generalization Under Data Imbalance
We begin by assessing whether models trained on imbalanced multilingual mixtures are still able to generalize across languages. Figure 2 reports the validation cross-entropy losses for models trained with varying proportions of English tokens, including extreme settings where English accounts for $90\%$ of the training data. Despite this imbalance, the models achieve competitive validation performance not only in English but also in minority languages such as Arabic, indicating that the learned representations capture transferable features rather than being restricted to high-resource languages. Together, these results suggest that multilingual circuits remain robust under imbalance and motivate a deeper investigation into how such robustness is organized within the model. More results are presented in Appendix E.
@@ 4.1.2 Language Entropy Across Layers
To investigate how multilingual representations are distributed within the network, we analyze the language entropy of feature activations at each layer. For each feature, we compute the multilingual score $H(f)$ as defined in Section 3.4, and multiply it by its corresponding activation rate computed over a large subset of the training data. Figure 3 shows the resulting average weighted entropy for the top100 score across layers for models trained on different data mixtures.
@@
The trend is robust across data mixtures and model scales. The emergence of a shared multilingual latent space in the middle layers is stable across different training data, and we also observe this well-defined behavior in the Llama 1B CLTs (see Figure 15), showing generalization to larger models.
@@
Model depth influences the behavior. For the 4-layer TinyStories model, the up-and-down pattern is absent and all layers have similar entropy scores, suggesting a minimum model size for this behavior to emerge.
The behavior is similar for the general entropy score. A discussion comparing the top100 score and the overall training distribution score is provided in Appendix D. These findings complement the validation loss results and provide quantitative evidence that the model organizes its internal representations to support cross-lingual generalization.
@@ 4.1.3 Multilingual Circuits Across Data Mixtures
To complement the multilingual score analysis, we extract attribution graphs to identify circuits active across data mixtures. We focus on two sentence types: (i) preposition sentences, where the model predicts a function word such as in J’ai bu une tasse or It was a piece, and (ii) content word sentences, where it predicts a meaningful token such as I prefer drinking tea to drinking in the 90% English mixture or Winter, spring, summer, and autumn are the four in the 20% mixture. We also analyze examples involving calendar terms (Monday, Tuesday, Wednesday, Thursday) and analogy prompts (the opposite of ’men’ is…). Figure 4 shows two representative circuits, one for an English preposition and one for a German content word, with remaining examples in Appendix F.
@@
Performance and representation quality are affected by tokenization. The model’s weaker performance on Arabic partially results from tokenization. Even with a balanced multilingual tokenizer, Arabic words are frequently split into small fragments, forcing the first layers (up to layer 3) to focus on reassembling words rather than learning higher-level meaning (Figures 21(b) and 22(b)).
Finding 4.
Feature alignment analysis shows no systematic bias toward English representations (Appendix L).

### v2

@@ Abstract
Multilingual Large Language Models (LLMs) can process many languages, yet how they internally represent this diversity remains unclear. Do they form shared multilingual representations with language-specific decoding, and if so, why does performance favor the dominant training language? To address this, we train models on different multilingual mixtures and analyze their internal mechanisms using Cross-Layer Transcoders (CLT) and Attribution Graphs. Our results reveal multilingual shared representations: the model employs highly similar features across languages, while language-specific decoding emerges in later layers.
Training models without English shows identical multilingual shared space structures. Decoding relies partly on a small set of high-frequency features in the final layers, which linearly encode language identity from early layers. Intervening on these features allows one language to be suppressed and another substituted. Finally, to explain non-English failures, we perform a Model-Diffing experiment: underperformance arises from dim late-layer features, weak middle-layer clusters, and tokenizer bias toward English that forces early layers to specialize in word reassembly. Finetuning strengthens these features and their links, improving token assembly and language-specific decoding, providing a mechanistic explanation for multilingual gaps.
Machine Learning, ICML
Figure 1: Cross-Layer Transcoders (CLTs) (Ameisen et al., 2025) (bottom) reveal that multilingual LLMs (middle) consistently form a shared language-agnostic space in middle layers (gradient middle region) where all languages converge, regardless of English dominance in the training mixtures (top).
@@ 1 Introduction
The inclusion of diverse languages in the large-scale training data of LLMs (Grattafiori et al., 2024; Team et al., 2024; Brown et al., 2020; Chowdhery et al., 2022) has led to remarkable multilingual capabilities (Shi et al., 2022; Conneau et al., 2020a; Workshop et al., 2023) However, the underlying mechanisms driving their multilingual behavior remain poorly understood. For example, do LLMs form shared multilingual representations in their internal layers, so, why does performance still favor the dominant training language, such as English? Prior mechanistic investigations have yielded conflicting findings about multilingual processing. While some work maps inputs into English-aligned latent spaces and suggests that LLMs process non-English inputs through English-biased representations (Wendler et al., 2024; Schut et al., 2025), other studies find evidence of shared cross-lingual grammatical structures encoded in common feature directions (Brinkmann et al., 2025). Moreover, all existing approaches rely primarily on static probing methods that cannot capture the dynamic transformation of linguistic information across model layers.
To address the aforementioned limitations, we derive attribution graphs from our self-trained cross-layer transcoders (Ameisen et al., 2025), a recent advance in mechanistic interpretability, which can track feature interactions across layers and provide the first mechanistic account of how LLMs transition between shared and language-specific processing modes. Our results suggest that all languages have a shared representation in LLMs’ internal layers, which we call as a “pivot language.” We also discover language-specific decoding in later layers to handle their multilingual output generation.
Depending on our CLT-driven insights, we conduct two causal studies. First, we are able to change the output language by intervening on a small set of high-frequency language features in the final layers decoding from the pivot language to the target language. Further, we investigate the causal effect of training data composition by varying English-to-multilingual ratios. We analyze our re-trained models on different language ratios, and reveal how language distribution during pretraining shapes internal multilingual representations. Our contributions are as follows:
@@
    We find that multilingual LLMs form shared multilingual representations, relying on consistent internal circuits to process equivalent queries across languages. These structures emerge even without English or a dominant language, indicating that pivot mechanisms are an architectural property rather than a consequence of language dominance. While English-dominant models may still preferentially align with English, it is not strictly required for the emergence of shared representations.
@@
    We release nine multilingual models [^1] Models: Multilingual TinyStories and Multilingual GPT-2 (two sizes × four data mixtures + non-English model), three finetuned models and their corresponding cross-layer transcoders (CLTs) [^2] CLTs: Multilingual CLTs, trained on balanced language distributions to support cross-lingual analysis and transfer. We also release our codebase [^3] Code: https://github.com/abirharrasse/MultilingualCLTs.
@@ 2.1 Multilingual Language Models
Multilingual language models have evolved from early architectures like mBERT (Devlin et al., 2019) to sophisticated models such as XLM-R (Conneau et al., 2020b), LLaMA (Touvron et al., 2023), and Gemma (Team et al., 2024). Despite their impressive capabilities, systematic performance disparities persist across languages (Pires et al., 2019; Hu et al., 2020). English consistently outperforms other languages, particularly for low-resource languages and complex reasoning tasks (Lauscher et al., 2020; Shi et al., 2022). Benchmarks like XTREME (Hu et al., 2020) and XNLI (Conneau et al., 2020b) have documented these gaps across language families. The English-dominant training data composition has been hypothesized as a primary factor (Bender et al., 2021), though the underlying mechanisms driving these disparities remain poorly understood.
@@ 2.2 Internal Representations and the Pivot Hypothesis
Recent investigations have revealed that multilingual LLMs process non-English inputs through English-aligned latent representations. Wendler et al. (2024) used logit lens analysis to show that intermediate layers predict English tokens before target-language tokens. Similarly, Schut et al. (2025) demonstrated systematic bias toward English-like representations regardless of input language. These findings suggest models internally translate to English, perform reasoning, then translate back, a mechanism that has been leveraged through embedding alignment (Lample & Conneau, 2019; Artetxe et al., 2018), translation-based fine-tuning (Zhu et al., 2024), and romanization strategies for non-Latin scripts (Saji et al., 2025).
Prompting strategies also exploit this English-pivot behavior, including direct prompt translation (Shi et al., 2022; Ahuja et al., 2023; Etxaniz et al., 2023) and English chain-of-thought reasoning (Wei et al., 2022; Vatsal et al., 2025). While some grammatical features exhibit universal representations across languages (Brinkmann et al., 2025), lexical and semantic processing appears more language-specific (Chi et al., 2020).
These observations align with the recently proposed “Platonic hypothesis” (Huh et al., 2024), which suggests that models converge toward abstract, unified representations capturing conceptual essences beyond surface forms. The shared multilingual space behavior, where models develop language-agnostic representations before language-specific decoding, provides some empirical support for this hypothesis in multilingual settings.
@@ 2.3 Mechanistic Interpretability in LLMs
Mechanistic interpretability seeks to understand transformer internals through circuit analysis (Olah et al., 2020; Elhage et al., 2021). Early work identified specific circuits for behaviors like indirect object identification (Wang et al., 2022) and induction heads (Olsson et al., 2022). Sparse Autoencoders (SAEs) decompose representations into interpretable features (Cunningham et al., 2023; Templeton et al., 2024), while transcoders directly model MLP outputs rather than autoencoding activations (Dunefsky et al., 2024).
Cross-Layer Transcoders (CLTs) advance this approach by assigning distinct decoder matrices per downstream layer, significantly simplifying feature graphs and enabling tractable circuit analysis (Ameisen et al., 2025; Hanna et al., 2025). However, CLT applications remain limited—primarily Anthropic’s foundational work (Lindsey et al., 2025b), a “greater-than” mechanism study (Merullo et al., 2025), and recent open-source extensions (Lindsey et al., 2025a). This scarcity reflects the considerable challenges in training CLTs at scale.
For multilingual models, mechanistic analyses are particularly sparse. Beyond Wendler et al. (2024)’s logit lens study and Anthropic’s case study (Lindsey et al., 2025b), our work represents the first comprehensive CLT-based investigation of multilingual processing. Crucially, ensuring fair cross-language comparisons requires training CLTs on balanced multilingual distributions to avoid feature bias toward dominant languages.
@@ 3.1 Mixture of multilingual LLMs
We study the multilingual behavior of LLMs across five languages: English, German, French, Arabic, and Chinese. To investigate how training data composition affects internal representations, we train four GPT-2 style models (12 layers, $\sim$177.6M parameters) on a realistic web mixture (OpenWebText + FineWeb2 (Penedo et al., 2025)), and four tiny-stories style models (4 layers and $\sim$68.5M parameters) on a controlled narrative dataset (TinyStories translated into all five languages). Finally, we include an off-the-shelf LLaMA-3.2-1B model (Meta AI, 2026) for comparison.
@@ 3.2 Cross-Layer Transcoders (CLTs)
To probe internal multilingual mechanisms, we employ CLTs, which map activations between layers to reveal how semantic representations evolve during processing. A CLT consists of neurons (features) divided into encoder and decoder components.
@@
Following Anthropic guidelines (Ameisen et al., 2025), the final training objective is
@@ 3.3 Graph attribution
$a_{\ell,k,n}^{\ell^{\prime},k^{\prime},n^{\prime}}=\sum_{\ell\leq s\leq\ell^{\prime}}f_{k,n}^{\ell\to s}\;J_{s,k}^{\ell^{\prime},k^{\prime}}\;g_{k^{\prime},n^{\prime}}^{\ell^{\prime}},$ (4)
where $f_{k,n}^{\ell\to s}$ denotes the vector for feature $n$ in the decoder matrix projecting from layer $\ell$ to $s$, $J_{s,k}^{\ell^{\prime},k^{\prime}}$ is the Jacobian between the MLP output at $(\ell,k)$ and the MLP input at $(\ell^{\prime},k^{\prime})$, computed during a forward pass where nonlinearities (normalization layers, attention computations, and MLPs) are frozen using stop-gradient operations, and $g_{k^{\prime},n^{\prime}}^{\ell^{\prime}}$ is the corresponding encoder feature at layer $\ell^{\prime}$ and position $k^{\prime}$. The graph is then pruned to retain only features that cumulatively account for $80\%$ of the effect on the final logit, and edges that cumulatively account for $95\%$ of the edge effect on the final logit. We use the circuit-tracer library (Hanna et al., 2025).
@@ 3.4 Multilingual Score
For each CLT feature $f$, we compute the multilingual scores on both the top 100 most activated sequences for that feature and on a general sample of 600k sequences balanced across the five languages (each of length 16). If a feature is active for at least one token in a sequence, it is considered active for that sequence. Since each sequence corresponds to a specific language, this allows us to compute a normalized activation distribution:
@@
which measures multilinguality: low $H(f)$ indicates language-specific features, while high $H(f)$ indicates multilingual features. We normalize the entropy scores to have scores between 0 and 1.
@@ 4 Results
    Do models form shared multilingual representations or a pivot language across data mixtures? (Section 4.1)
@@
    How do models decode language identity and route information to the correct output language? (Section 4.2)
@@
    Why do models still fail in non-English settings despite shared representations? (Section 4.3)
@@ 4.1.1 Generalization Under Data Imbalance
We begin by assessing whether models trained on imbalanced multilingual mixtures are still able to generalize across languages. Figure 11 reports the validation cross-entropy losses for models trained with varying proportions of English tokens, including extreme settings where English accounts for $90\%$ of the training data. Despite this imbalance, the models achieve competitive validation performance not only in English but also in minority languages such as Arabic, indicating that the learned representations capture transferable features rather than being restricted to high-resource languages. Together, these results suggest that multilingual circuits remain robust under imbalance and motivate a deeper investigation into how such robustness is organized within the model. More results are presented in Appendix C.
@@ 4.1.2 Language Entropy Across Layers
To investigate how multilingual representations are distributed within the network, we analyze the language entropy of feature activations at each layer. For each feature, we compute the multilingual score $H(f)$ as defined in Section 3.4, and multiply it by its corresponding activation rate computed over a large subset of the training data (4M Tokens). Figure 2 shows the resulting average weighted entropy for the top100 score across layers for models trained on different data mixtures.
@@
The trend is robust across data mixtures and model scales. The emergence of a shared multilingual latent space in the middle layers is stable across different training data, and we also observe this well-defined behavior in the Llama-3.2-1B CLTs (see Figure 20), showing generalization to larger models.
@@
Model depth influences the behavior. For the 4-layer TinyStories model, the up-and-down pattern is absent and all layers have similar entropy scores (Figure 20), suggesting a minimum model size for this behavior to emerge.
The behavior is similar for the general entropy score. A discussion comparing the top100 score and the overall training distribution score is provided in Appendix E. These findings complement the validation loss results and provide quantitative evidence that the model organizes its internal representations to support cross-lingual generalization.
Shared Multilingual Space Independence from English.
To validate that shared multilingual space structures do not require English dominance, we trained GPT-2 on balanced non-English data (25% French, 25% German, 25% Arabic, 25% Chinese) and computed weighted multilingual entropy across layers. The resulting curve exhibits the identical characteristic rise-and-fall pattern observed in English-dominant models, with peak entropy in layers 5-6 (Figure 22). This demonstrates that the shared multilingual space emergence is a fundamental architectural property, not an artifact of English dominance.
@@ 4.1.3 Multilingual Circuits Across Data Mixtures
To complement the multilingual score analysis, we extract attribution graphs to identify circuits active across data mixtures. We focus on two sentence types: (i) preposition sentences, where the model predicts a function word such as in J’ai bu une tasse or It was a piece, and (ii) content word sentences, where it predicts a meaningful token such as I prefer drinking tea to drinking in the 90% English mixture or Winter, spring, summer, and autumn are the four in the 20% mixture. We also analyze examples involving calendar terms (Monday, Tuesday, Wednesday, Thursday) and analogy prompts (the opposite of ’men’ is…). Figure 3 shows two representative circuits, one for an English preposition and one for a German content word, with remaining examples in Appendix G.1 and more details about the clustering and cluster’s entropy calculation in Appendix G .
@@
Performance and representation quality are affected by tokenization. The model’s weaker performance on Arabic partially results from tokenization. Even with a balanced multilingual tokenizer, Arabic words are frequently split into small fragments, forcing the first layers (up to layer 3) to focus on reassembling words rather than learning higher-level meaning (Figures 25(b) and 26(b)).